# Seeing What’s Not There: Spurious Correlation in Multimodal LLMs

### v1

@@ Appendix I SpurLens Validation
81 ptarmigan grass 2.0 rock 5.3 lichen 32.0 ice 4.0
82 ruffed grouse bush 12.0 branch 37.7 grass 14.0 branch 26.7
89 sulphur-crested cockatoo seed 0.7 tree 29.3 seed 12.0 seed 4.7
91 coucal bark 4.0 bark 18.0 tree 28.7 sunlight 8.0
94 hummingbird nectar 3.3 nectar 6.0 nectar 8.0 nectar 3.3
103 platypus sand 4.0 leaf 9.3 branch 18.0 bird 6.7
105 koala foliage 0.0 foliage 6.0 foliage 2.0 foliage 2.0
147 grey whale coral 2.0 fisherman 36.0 fisherman 6.7 beach 16.7
288 leopard thicket 0.0 sky 16.7 antelope 4.7 tree 18.7
@@
309 bee pollen 2.0 sky 2.0 pollen 2.7 sky 2.0
313 stick insect spider web 4.0 grass blades 10.0 bark 10.7 bark 4.0
324 small white butterfly grass 3.3 petal 32.7 petal 12.0 trail 8.7
325 sulphur butterfly stone 6.0 fruit 6.7 flower 3.3 fence 6.7
@@
351 hartebeest path 2.0 cloud 16.0 plain 38.0 plain 6.7
364 three-toed sloth rain 2.0 rock 13.3 tree 12.0 rain 4.7
@@
384 indri tree 12.3 tree 20.0 branch 21.3 tree 60.7
389 snoek fish rod 3.3 cooler 14.7 driftwood 8.0 tide pool 21.3
394 sturgeon bubble 3.5 bait 9.3 weed 12.7 riptide 6.7
395 gar fish sand 9.3 log 28.7 aquatic plants 6.0 aquatic plants 40.0
400 academic gown globe 2.7 diploma 5.3 diploma 4.0 flower 6.0
@@
419 Band-Aid bandaid box 11.3 alcohol wipes 18.0 skin 12.0 bandaid box 13.3
@@
425 barn fence 2.0 haystack 6.0 windmill 6.0 windmill 9.3
433 swimming cap swim goggles 4.0 paddle 13.3 swimwear 9.3 water bottle 8.0
434 bath towel sink 2.7 mirror 33.3 shower curtain 2.0 shower curtain 9.3
435 bathtub brush 4.0 brush 14.0 sink 53.3 wall tiles 10.0
440 beer bottle table 2.0 table 12.0 chair 6.0 table 5.3
@@
454 bookstore picture frame 5.3 picture frame 14.0 picture frame 2.7 picture frame 4.7
465 bulletproof vest plant 4.0 tactical belt 33.3 tactical belt 23.3 smoke grenade 8.7
466 high-speed train bridge 4.0 bridge 7.3 traffic light 15.3 bridge 4.7
490 chain mail tent 4.7 stone wall 14.0 dirt road 7.3 forest 15.3
@@
515 cowboy hat saddle 3.3 mountain 13.3 horse 4.0 fence 8.0
516 cradle night light 5.3 nursery rug 24.0 nursery rug 11.3 photo frame 7.3
525 dam crane 4.0 building 16.0 bicycle 4.0 crane 4.7
537 dog sled trail markers 2.7 ski tracks 28.0 ski tracks 21.3 ski tracks 23.3
543 dumbbell exercise ball 0.7 fitness tracker 12.7 floor tiles 8.0 fitness tracker 2.7
554 fireboat ladder 5.3 skyline 15.3 ladder 30.0 ladder 12.3
557 flagpole cloud 4.0 flower 7.3 cloud 8.0 sign 6.7
563 fountain pen stapler 1.9 chair 6.7 paper 1.3 stamp 11.1
565 freight car crane 0.9 railroad sign 4.0 crane 4.0 billboard -0.5
576 gondola streetlamps 2.0 paddle 9.3 paddle 5.3 paddle 6.8
585 hair spray conditioner 12.6 makeup 34.0 makeup 24.7 conditioner 18.7
@@
592 hard disk drive desk -1.3 keyboard 4.0 printer 4.7 book -1.3
602 gymnastic horizontal bar footwear 6.0 banner 9.3 footwear 4.0 footwear 5.3
626 lighter coaster 2.0 candle 12.0 cigarette 16.7 matchstick 6.7
@@
667 graduation cap diploma 4.0 diploma 5.3 gown 4.7 diploma 4.0
671 mountain bike backpack 4.7 trail 12.7 bush 8.0 backpack 15.3
678 neck brace lamp 11.3 wall 26.7 artwork 9.3 rug 6.7
680 baby pacifier blanket 46.0 blanket 48.0 swaddle 30.0 stroller 42.0
682 obelisk cloud 2.0 mountain 4.0 cloud 2.0 cloud 3.3
684 ocarina headphone 2.7 rug 2.0 cushion 12.0 headphone 5.3
695 padlock table 2.0 post 6.0 gate 2.7 notebook 2.0
702 parallel bars gym mat 4.0 gym bag 12.7 yoga mat 15.3 spotting belt 10.0
709 pencil case globe 4.0 paper 30.0 notebook 50.0 desk 1.3
720 pill bottle bandage 3.3 thermometer 4.7 thermometer 6.0 bandage 4.7
722 ping-pong ball referee 6.7 racket 25.3 player 37.3 floor 8.0
728 plastic bag fruit peels 2.7 water bottle 8.7 sidewalk 3.3 flower pot 6.7
731 plunger sponge 13.3 faucet 22.0 bathroom 44.7 trashcan 12.7
733 pole road 3.3 cloud 12.0 flag 10.7 cloud 8.0
737 soda bottle glass 4.7 chair 20.0 chair 14.0 burger 7.3
738 plant pot fertilizer 8.7 chair 12.0 chair 9.3 basket 8.0
739 potter’s wheel display 2.7 brush 21.3 clay 14.9 glaze 13.3
746 hockey puck skate 9.3 skate 13.3 player equipment 20.7 ice shavings 18.7
@@
766 rotisserie chair 13.3 rug 14.0 window 14.7 plant 7.3
767 eraser desk 5.3 notebook 14.0 binder 10.7 pen 4.0
@@
776 saxophone drink 4.0 fruit bowl 4.7 photo 20.7 microphone 14.0
785 seat belt driver 2.0 headrest 7.3 rearview mirror 6.0 cup holder 4.0
@@
792 shovel soil 4.3 wheelbarrow 6.7 rock 8.0 dirt 8.0
793 shower cap towel 4.0 towel 19.3 mirror 6.7 towel 4.7
797 sleeping bag hiking boots 6.7 hiking boots 18.7 trekking poles 14.7 map 12.0
801 snorkel sand 9.3 beach 14.0 wetsuit 8.0 beach 23.3
802 snowmobile trail markers 2.0 snow 12.7 snow 68.0 tree 4.0
803 snowplow snow 3.3 snow 12.7 snow 22.0 snow 14.0
@@
827 stove spice 10.0 oven 23.3 oven 50.7 pan 16.7
867 semi-trailer truck concrete barrier 4.0 pedestrian 12.7 pedestrian 6.0 concrete barrier 14.0
908 airplane wing propeller 3.3 propeller 6.0 sunset 4.0 sky 7.3
@@
0 tench net 6.0 fishing gear 2.0 aquatic plants 26.7 net 2.0
2 great white shark swimmer 3.3 buoy 0.0 buoy 0.0 buoy 0.0
14 indigo bunting grass 4.7 grass 0.0 branch 9.3 grass 0.0
42 agama fence 1.3 grass 0.7 grass 31.3 grass 0.0
50 American alligator water 1.3 fishing gear 1.3 fishing gear 4.7 water 4.0
58 water snake reed 4.0 reed 2.0 log 4.0 shell 1.3
@@
81 ptarmigan heath 26.0 heath 8.7 snow 19.3 heath 10.7
@@
89 sulphur-crested cockatoo tree 2.0 play gym 1.3 tree 22.7 feeder 0.0
@@
94 hummingbird bloom 1.3 sky 0.7 petal 32.7 feeder 0.0
103 platypus water 2.0 sand 0.7 lily pads 1.3 lily pads 0.7
105 koala waterhole 0.0 grass 0.0 foliage 6.0 foliage 2.0
147 grey whale wave 17.3 wave 4.7 wave 10.0 wave 2.7
288 leopard thicket 0.0 thicket 0.0 tree 6.0 fallen leaves 0.7
297 sloth bear tree 14.0 fungi 0.7 tree 45.3 sunlight 0.7
309 bee leaf 2.0 wood 3.3 flower 5.3 leaf 2.0
313 stick insect water droplet 6.0 leaf 0.7 flower petals 4.7 moss 2.0
324 small white butterfly wildflower 11.3 sun 4.7 wildflower 14.7 sun 2.0
325 sulphur butterfly petal 22.0 petal 6.0 petal 36.7 petal 2.0
335 fox squirrel acorn 2.7 flowerbed 0.7 tree 16.0 bird feeder 0.0
351 hartebeest fence 2.0 cactus 2.0 grass 6.7 muddy ground 0.0
364 three-toed sloth tree 4.0 tree 0.7 tree 25.3 rain 0.7
@@
384 indri tree 2.0 sunlight 0.7 tree 41.3 tree 8.0
389 snoek fish sea surface 21.3 sea surface 2.0 water 19.3 reel 2.7
@@
395 gar fish buoy 20.0 water surface 6.7 water surface 59.3 water surface 1.3
400 academic gown book 6.0 book 5.3 book 2.0 book 2.0
415 bakery parking lot 4.0 parking lot 11.3 streetlight 5.3 streetlight 3.3
@@
419 Band-Aid skin 7.3 skin 4.0 skin 6.7 skin 2.7
424 barbershop bench 2.0 bench 3.3 fence 0.0 fire hydrant 0.7
425 barn fence 2.0 fence 10.0 fence 9.3 fence 2.7
433 swimming cap towel 10.0 swimwear 2.0 snorkel 32.0 change rooms 4.7
@@
465 bulletproof vest backpack 2.0 boot 1.3 helmet 22.7 map 0.0
466 high-speed train landscape 0.0 landscape 0.0 landscape 0.0 landscape 0.0
490 chain mail shield 4.7 castle 1.3 stone wall 5.3 shield 2.0
491 chainsaw work boots 4.7 bark 1.3 workbench 12.0 work boots 2.0
515 cowboy hat boot 6.7 chap 4.7 boot 13.3 chap 3.3
516 cradle curtain 1.3 stroller 10.0 stroller 6.0 stroller 1.3
@@
537 dog sled ski tracks 13.3 pine trees 0.7 pine trees 2.7 horizon 0.0
543 dumbbell suspension straps 3.3 floor tiles 2.0 gym mat 2.0 floor tiles 2.0
@@
565 freight car road 6.0 road 8.7 road 22.0 pipe 0.0
576 gondola bridge 8.0 bridge 3.3 cafés 24.0 streetlamps 0.0
585 hair spray mirror 25.3 mirror 10.7 mirror 27.3 sink 5.3
588 hamper basket 2.7 basket 53.3 basket 30.7 basket 23.3
592 hard disk drive computer 4.0 laptop 8.0 keyboard 4.7 desk 2.0
602 gymnastic horizontal bar banner 11.3 banner 6.7 banner 12.0 banner 7.3
626 lighter bag 15.3 candle 14.0 candle 11.3 keychain 1.3
@@
667 graduation cap photo 4.0 photo 4.0 rope 10.7 photo 4.0
671 mountain bike water bottle 2.7 backpack 4.7 backpack 4.0 hill 2.0
678 neck brace cup 2.7 pillow 4.0 cup 2.0 coaster 1.3
680 baby pacifier crib 10.7 crib 3.3 crib 22.0 blanket 2.0
682 obelisk planter 15.3 planter 15.3 flagpole 53.3 cloud 2.0
684 ocarina guitar 3.3 sheet music 5.3 drum 12.7 drum 0.0
695 padlock wall 4.0 door 14.7 chain 25.3 drawer 6.0
@@
720 pill bottle tissue 14.7 tissue 11.3 glass 12.7 clipboard 1.3
722 ping-pong ball racket 3.3 light 2.0 racket 14.7 racket 1.3
@@
731 plunger toilet 7.3 countertop 7.3 pipe 27.3 toilet 0.7
733 pole light 30.0 light 22.0 light 35.3 light 12.0
737 soda bottle glass 18.0 chair 18.0 glass 14.7 cup 13.3
738 plant pot window 14.7 window 20.7 window 13.3 gardening gloves 11.3
739 potter’s wheel tool 2.7 shelf 2.0 stool 8.0 display 0.0
746 hockey puck jersey 2.0 jersey 2.7 pant 0.7 zamboni 0.0
749 quill paper 0.7 candle 2.0 pen holder 14.7 paper 3.3
@@
755 radio telescope power lines 0.0 power lines 6.0 power lines 6.0 power lines 0.0
756 rain barrel bucket 11.3 potted plants 13.3 bucket 39.3 bucket 6.0
766 rotisserie cutlery 4.7 cutlery 4.7 cutlery 5.3 vegetable 0.0
@@
771 safe lamp 2.0 bookshelf 14.7 lamp 0.7 medal 0.7
776 saxophone guitar 6.0 guitar 2.7 guitar 4.7 lamp 0.0
785 seat belt headrest 15.3 rearview mirror 8.0 driver 8.7 safety light 0.0
788 shoe store parked car 9.3 clock 3.3 poster 2.0 vendor cart 2.7
@@
793 shower cap bathrobe 15.3 mirror 10.7 towel 62.7 bathrobe 3.3
797 sleeping bag blanket 13.3 blanket 3.3 blanket 22.0 backpack 4.7
@@
802 snowmobile winter jacket 4.7 winter jacket 2.7 winter jacket 6.7 winter jacket 1.3
803 snowplow shovel 8.7 road 0.7 shovel 8.7 shovel 3.3
@@
827 stove oven 36.7 oven 16.0 oven 22.7 oven 18.0
867 semi-trailer truck highway 10.0 parking lot 20.7 highway 38.0 guardrail 2.0
908 airplane wing star 2.0 star 2.7 tree 0.7 building 0.0
933 cheeseburger plate 0.0 plate 0.0 plate 0.0 plate 0.0

### v2

@@ Appendix I Hyperparameter Sensitivity Analysis
81 ptarmigan sky 2.0 boulder 4.0 mountain 25.3 sky 2.0
82 ruffed grouse bush 12.0 branch 34.7 grass 14.0 branch 26.7
89 sulphur-crested cockatoo feeder 0.0 tree 29.3 feeder 8.0 bush 4.0
91 coucal nest 10.0 bark 18.0 tree 28.7 nest 12.0
94 hummingbird sky 2.0 sky 4.0 feeder 6.7 sky 2.0
103 platypus bubble 4.7 bubble 23.3 branch 18.0 bubble 9.3
105 koala eucalyptus leaves 1.3 foliage 6.0 foliage 2.0 eucalyptus leaves 4.0
147 grey whale dolphin 4.0 fisherman 36.0 dolphin 29.3 beach 16.7
288 leopard fence 0.0 tree 16.0 waterhole 8.0 tree 18.7
@@
309 bee grass 0.0 garden 6.0 garden 0.0 garden 0.7
313 stick insect stone 1.3 water droplet -1.3 dirt mound 1.3 dirt mound 3.3
324 small white butterfly grass 3.3 garden 19.3 garden 14.7 wildflower 3.3
325 sulphur butterfly stone 6.0 twig 6.7 bee 8.0 grass 5.3
@@
351 hartebeest muddy ground 0.0 rock 15.3 waterhole 18.7 horizon 5.3
364 three-toed sloth bird 2.0 rock 13.3 tree 12.0 bird 2.0
@@
384 indri tree 12.0 tree 20.0 branch 21.3 tree 60.7
389 snoek fish bait 7.3 cooler 14.7 bait 24.0 tide pool 21.3
394 sturgeon camera 1.7 glove 5.3 riverbed 20.7 net 5.3
395 gar fish bait 12.0 bait 27.3 bait 13.3 turtle 29.3
400 academic gown desk 2.0 diploma 5.3 diploma 4.0 flower 6.0
@@
419 Band-Aid alcohol wipes 0.7 alcohol wipes 18.0 skin 12.0 skin 12.7
@@
425 barn bucket 2.0 haystack 6.0 windmill 6.0 windmill 9.3
433 swimming cap bleacher 4.7 paddle 13.3 lifeguard 17.3 water bottle 8.0
434 bath towel sink 2.7 mirror 33.3 shelf 2.0 showerhead 5.3
435 bathtub brush 4.0 brush 14.0 soap 48.0 wall tiles 10.0
440 beer bottle chip 2.0 table 12.0 chair 6.0 table 5.3
@@
454 bookstore picture frame 5.3 picture frame 14.0 laptop 3.3 laptop 7.3
465 bulletproof vest picture 0.0 tactical belt 33.3 tactical belt 23.3 tactical belt 7.3
466 high-speed train overpass 9.3 overpass 14.0 overpass 22.7 overpass 12.0
490 chain mail tent 4.7 stone wall 14.0 sword stand 7.3 forest 15.3
@@
515 cowboy hat saddle 3.3 guitar 10.7 plain 4.0 fence 8.0
516 cradle changing table 6.0 nursery rug 24.0 nursery rug 11.3 photo frame 7.3
525 dam bridge 13.3 bridge 36.0 bridge 14.0 bridge 14.7
537 dog sled ski tracks 2.7 ski tracks 28.0 ski tracks 21.3 ski tracks 23.3
543 dumbbell kettlebell 6.0 fitness tracker 12.7 floor tiles 8.0 yoga block 6.0
554 fireboat tugboat 24.7 coast guard 29.3 tugboat 46.0 tugboat 35.7
557 flagpole cloud 4.0 bench 6.0 cloud 8.0 sign 6.7
563 fountain pen stapler 2.0 ink bottle 8.0 ink bottle 8.7 stamp 10.0
565 freight car crane 0.7 warehouse 7.3 crane 4.0 warehouse 3.3
576 gondola flag 2.0 canal 12.7 paddle 5.3 canal 7.3
585 hair spray bobby pins 14.3 bobby pins 38.0 bobby pins 16.7 conditioner 18.7
@@
592 hard disk drive desk -1.3 keyboard 4.0 printer 4.7 wall -2.7
602 gymnastic horizontal bar exercise balls 8.7 ring 9.3 jump rope 7.3 ring 7.3
626 lighter cell phone 0.7 cell phone 7.3 keychain 16.7 matchstick 6.7
@@
667 graduation cap flag 4.0 diploma 5.3 gown 4.7 diploma 4.0
671 mountain bike backpack 4.7 rock 12.7 bush 8.0 backpack 15.3
678 neck brace lamp 11.3 wall 26.7 magazine 8.0 lamp 6.0
680 baby pacifier teething toy 72.0 teething toy 75.3 teething toy 40.0 teething toy 72.7
682 obelisk sidewalk 2.0 mountain 4.0 sidewalk 2.0 cloud 3.3
684 ocarina headphone 2.7 photo -1.3 table 0.0 headphone 5.3
695 padlock bag 2.0 post 6.0 box 4.7 gate 2.0
702 parallel bars pull-up bar 9.3 wrist tape 15.3 pull-up bar 20.0 pull-up bar 10.7
709 pencil case clip 12.0 clip 36.0 clip 58.7 clip 8.7
720 pill bottle first aid kit 4.7 inhaler 19.3 glass of water 11.3 inhaler 6.7
722 ping-pong ball referee 7.3 racket 25.3 racket 36.0 floor 8.0
728 plastic bag tissue 2.0 water bottle 8.7 tissue 3.3 tissue 2.7
731 plunger soap 12.7 faucet 22.0 tile 42.0 trashcan 12.7
733 pole road 3.3 cloud 12.0 cloud 9.3 cloud 8.0
737 soda bottle table 2.7 chair 20.0 chair 14.0 straw 4.7
738 plant pot vase 16.7 vase 15.3 vase 19.3 vase 12.0
739 potter’s wheel display 2.7 brush 21.3 clay 12.6 water 13.3
746 hockey puck helmet 9.3 skate 13.3 skate 18.7 game clock 16.7
@@
766 rotisserie bread rolls 11.3 food baskets 19.3 window 14.7 bread rolls 18.0
767 eraser marker 9.3 highlighter 16.7 sticky note 24.0 marker 6.0
@@
776 saxophone bass guitar 4.0 music stand 4.0 photo 20.7 microphone 14.0
785 seat belt driver 2.0 headrest 7.3 rearview mirror 6.0 driver 2.7
@@
792 shovel soil 4.7 wheelbarrow 6.7 rock 8.0 garden tool 6.7
793 shower cap comb 6.0 towel 19.3 mirror 6.7 comb 8.7
797 sleeping bag tent 7.3 hiking boots 18.7 campsite 17.3 ground mat 14.7
801 snorkel sand 9.3 beach 14.0 goggles 19.3 beach 23.3
802 snowmobile trail markers 2.0 snow shovel 16.0 snow 68.0 snow shovel 8.0
803 snowplow utility pole 2.0 road 8.0 snowdrift 21.3 snowdrift 6.0
@@
827 stove spoon 9.3 pan 20.0 sink 19.3 pan 16.7
867 semi-trailer truck traffic light 4.0 pedestrian 12.7 pedestrian 6.0 concrete barrier 14.0
908 airplane wing propeller 3.3 propeller 6.0 sea 2.0 sky 7.3
@@
0 tench net 6.0 lure 2.7 lure 20.7 net 2.0
2 great white shark swimmer 3.3 net 0.0 net 0.0 net 0.0
14 indigo bunting feeder 7.3 branch 0.0 feeder 11.3 branch 0.0
42 agama mud 0.7 tree 0.7 grass 31.3 stone 0.0
50 American alligator water 1.3 tree 0.0 water 0.7 water 4.0
58 water snake reed 4.0 reed 2.0 shell 3.3 shell 1.3
@@
81 ptarmigan grass 24.0 snowdrift 6.0 grass 14.7 grass 9.3
@@
89 sulphur-crested cockatoo tree 2.0 feeder 0.0 tree 22.7 feeder 0.0
@@
94 hummingbird garden 0.7 sky 0.7 petal 32.7 tree 0.0
103 platypus frog 2.0 sand 0.7 lily pads 1.3 lily pads 0.7
105 koala branch 0.0 branch 0.0 foliage 6.0 foliage 2.0
147 grey whale fisherman 14.7 surfboard 0.7 fisherman 4.0 surfboard 2.7
288 leopard fence 0.0 fence 0.0 tree 6.0 fence 0.0
297 sloth bear tree 14.0 tree 0.7 tree 45.3 waterhole 0.0
309 bee leaf 2.0 flower 0.7 flower 5.3 leaf 2.0
313 stick insect water droplet 6.0 dirt mound 0.0 dirt mound 0.0 dirt mound 0.0
324 small white butterfly wildflower 11.3 tree 1.3 wildflower 14.7 leaf 2.0
325 sulphur butterfly petal 22.0 petal 6.0 petal 36.7 garden 2.0
335 fox squirrel acorn 2.7 acorn 0.7 tree 16.0 tree 0.0
351 hartebeest muddy ground 0.0 tree 2.0 grass 6.7 muddy ground 0.0
364 three-toed sloth tree 4.0 tree 0.7 tree 25.3 tree 0.0
@@
384 indri tree 2.0 rock 0.0 tree 41.3 tree 8.0
389 snoek fish seaweed 24.0 sea surface 2.0 seaweed 30.0 rod 4.7
@@
395 gar fish turtle 26.0 turtle 11.3 turtle 50.0 bird 0.0
400 academic gown diploma 27.3 diploma 30.7 diploma 32.0 diploma 25.3
415 bakery car 2.7 car 5.3 car 3.3 car 1.3
@@
419 Band-Aid blood 9.3 alcohol wipes 6.7 skin 6.7 ruler 4.0
424 barbershop bus stop 4.0 bus stop 4.0 flowerpot 0.0 fire hydrant 0.7
425 barn fence 2.0 birdhouse 11.3 fence 9.3 fence 2.7
433 swimming cap towel 10.0 pool 2.0 snorkel 32.0 water bottle 2.7
@@
465 bulletproof vest gun 4.7 gun 2.0 helmet 22.7 floor 0.0
466 high-speed train construction barriers 0.0 construction barriers 0.0 construction barriers 0.0 construction barriers 0.0
490 chain mail sword stand 7.3 torch 4.7 sword stand 12.7 shield 2.0
491 chainsaw work boots 4.7 woodpile 1.3 workbench 12.0 work boots 2.0
515 cowboy hat horse 10.7 whip 7.3 guitar 16.7 whip 4.0
516 cradle bedside table 5.3 stroller 10.0 infant toys 6.7 baby blanket 4.0
@@
537 dog sled ski tracks 13.3 pine trees 0.7 pine trees 2.7 canoe 0.0
543 dumbbell weight plate 2.7 weight plate 2.0 weight plate 3.3 weight plate 2.0
@@
565 freight car road 6.0 road 8.7 road 22.0 grass 0.0
576 gondola bridge 8.0 bridge 3.3 poster 12.7 tree 0.0
585 hair spray shampoo 31.3 shampoo 17.3 shampoo 42.0 shampoo 5.3
588 hamper wall 1.3 curtain 15.3 wall 12.0 curtain 6.0
592 hard disk drive router 6.7 printer 12.7 router 7.3 desk 2.0
602 gymnastic horizontal bar bench 15.3 bench 10.7 bench 13.3 banner 7.3
626 lighter matchbox 26.0 cigarette 18.0 cigarette 25.3 keychain 1.3
@@
667 graduation cap diploma 31.3 diploma 30.0 diploma 36.0 diploma 30.0
671 mountain bike water bottle 2.7 backpack 4.7 backpack 4.0 glove 2.0
678 neck brace cup 2.7 remote 3.3 cup 2.0 wall 0.7
680 baby pacifier teddy bear 16.0 diaper 4.0 crib 22.0 crib 2.0
682 obelisk lamp post 11.3 statue 12.0 statue 20.0 pathway 2.0
684 ocarina photo 1.3 microphone 2.7 headphone 2.7 headphone 0.0
695 padlock wall 4.0 door 14.7 chain 25.3 wall 4.0
@@
720 pill bottle tissue 14.7 tissue 11.3 thermometer 9.3 coffee cup 0.0
722 ping-pong ball referee 4.7 scoreboard 2.7 racket 14.7 scoreboard 2.0
@@
731 plunger cleaner 11.3 brush 12.0 pipe 27.3 brush 1.3
733 pole street 54.0 building 52.0 car 46.7 street 52.0
737 soda bottle table 14.0 chair 18.0 cup 14.0 cup 13.3
738 plant pot watering can 18.7 window 20.7 watering can 34.7 watering can 23.3
739 potter’s wheel sculpture 2.7 shelf 2.0 jug 6.7 jug 0.0
746 hockey puck jersey 2.0 penalty box 2.7 skate 2.0 penalty box 1.3
749 quill notebook 0.7 candle 2.0 notebook 15.3 ink bottle 4.7
@@
755 radio telescope satellite 2.0 satellite 15.3 power lines 6.0 bench 0.0
756 rain barrel potted plants 8.0 shed roof 13.3 garden statue 19.3 potted plants 4.7
766 rotisserie cutlery 4.7 serving trays 4.7 cutlery 5.3 plant 0.0
@@
771 safe electronic devices 1.3 bookshelf 14.7 clock 0.0 window 0.0
776 saxophone guitar 6.0 guitar 2.7 guitar 4.7 amplifier 0.0
785 seat belt headrest 15.3 rearview mirror 8.0 driver 8.7 dashboard 0.0
788 shoe store vendor cart 7.3 clock 3.3 poster 2.0 vendor cart 2.7
@@
793 shower cap bathrobe 15.3 mirror 10.7 towel 62.7 towel 3.3
797 sleeping bag camp chair 14.7 camp chair 11.3 camp chair 16.0 backpack 4.7
@@
802 snowmobile winter jacket 4.7 trail markers 4.7 trail markers 7.3 winter jacket 1.3
803 snowplow shovel 8.7 glove 0.7 shovel 8.7 shovel 3.3
@@
827 stove pan 18.0 dishwasher 17.3 pan 14.7 cabinet 12.7
867 semi-trailer truck construction site 16.7 construction site 22.0 construction site 40.7 construction site 5.3
908 airplane wing propeller 5.3 propeller 4.7 propeller 3.3 propeller 1.3
933 cheeseburger napkin 0.0 burger wrapper 2.0 burger wrapper 2.0 napkin 0.0